OpenStudy (anonymous):

Write equation in standard form. Find the canter (h, k) and radius r for x^2 + y^2 + 4y - 12 = 0

OpenStudy (anonymous):

\[ x^2 + y^2 + 4y - 12 = 0\\ x^2 + y^2 +4 y + 4 -4 -12=0 \] Can you finish it?

OpenStudy (anonymous):

\[ x^2 + y^2 +4 y + 4 -4 -12=0\\ x^2 +(y+2)^2 =16=4^2 \]

OpenStudy (anonymous):

Are you still with me?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

yup I got it now.